# GUS-Net: Social Bias Classification in Text with Generalizations, Unfairness, and Stereotypes

### v2

@@ Abstract
The detection of bias in natural language processing (NLP) is a critical challenge, particularly with the increasing use of large language models (LLMs) in various domains. This paper introduces GUS-Net, an innovative approach to bias detection that focuses on three key types of biases: Generalizations, Unfairness, and Stereotypes. GUS-Net leverages generative AI and automated agents to create a comprehensive synthetic dataset, enabling robust multi-label token classification. Our methodology enhances traditional bias detection methods by incorporating the contextual embeddings of pre-trained models, resulting in improved accuracy and depth in identifying biased entities. Through extensive experiments, we demonstrate that GUS-Net outperforms state-of-the-art techniques, achieving superior performance in terms of accuracy, F1-score, and Hamming Loss. The findings highlight GUS-Net’s effectiveness in capturing a wide range of biases across diverse contexts, making it a valuable tool for social bias detection in text. This study contributes to the ongoing efforts in NLP to address implicit bias, providing a pathway for future research and applications in various fields. The Jupyter notebooks used to create the dataset and model are available at https://github.com/Ethical-Spectacle/fair-ly/tree/main/resources.
Warning: This paper contains examples of harmful language, and reader
discretion is recommended.
@@ 1 Introduction
The detection of bias in natural language processing (NLP) [13] is an important task, particularly with the increasing use of large language models (LLMs) [33] in domains like education [18] and business [29]. Bias can significantly influence public perception and decision-making, often subtly reinforcing stereotypes or propagating discriminatory practices. While explicit bias, which refers to clearly expressed prejudice or favoritism, is easy to define, implicit bias involves more subtle and often unconscious associations or attitudes. Therefore, identifying and mitigating implicit bias in the text is challenging: what is perceived as biased can vary greatly depending on the context, including the perspectives of viewers and speakers. For example, consider the phrase “hard-working immigrants”. To some, this phrase may appear positive, acknowledging the effort and diligence of immigrants. However, from another perspective, it might be perceived as implicitly biased, suggesting that immigrants are expected to work harder than others to be valued or accepted. This subtle implication can be seen as reinforcing a stereotype that separates immigrants from native citizens, placing an undue burden of proof on their worthiness. This subjectivity underlines the complexity of implicit bias detection, making it a critical area of research within NLP [14, 28, 25].
While the implicit nature of bias can manifest in subtle forms, such as the choice of words, framing of narratives, or the omission of certain viewpoints, traditional approaches to bias detection have typically relied on human annotators to label datasets [25, 26]. Although this method has been essential for creating foundational resources, it is susceptible to the biases of annotators, who may struggle to step outside their own ideological frameworks. As a result, existing datasets often lack the diversity of viewpoints necessary to capture implicit biases effectively. Additionally, many datasets, such as the one utilized by the Dbias model [32], are limited in scope, focusing on a narrow range of biases and failing to generalize across different contexts. The Nbias framework [25], while an advancement in incorporating named entity recognition (NER) tasks, still emphasizes explicit biases and overlooks the structural elements of implicit bias, such as generalizations and stereotypes. Despite studies of robust annotations conducted by trained experts [27], all the previous bias datasets rely on human annotators, which means that these datasets often lack the diversity of viewpoints necessary to capture implicit bias.
In response to these challenges, this paper introduces the Generalizations, Unfairness, and Stereotypes Network (GUS-Net), which stands for three kinds of biases that are commonly considered in legal and psychological literature [1, 3, 12, 21]: Generalizations, Unfairness, and Stereotypes. GUS-Net leverages generative AI and automated agents to build an optimal dataset for bias detection. Using the synthetic data generated by these automated agents, we further fine-tune the pre-trained model BERT (Bidirectional Encoder Representations from Transformers) [10] for the task of multi-label token classification. This approach improves upon traditional methods by combining LLM reasoning and the powerful contextual embeddings provided by pre-trained models, resulting in more accurate and comprehensive bias detection across various types of text. Our extensive experiments demonstrate that the proposed method outperforms state-of-the-art techniques in terms of dataset diversity and annotation depth. The fine-tuned model not only achieves superior performance on traditional metrics, such as accuracy and F1-score but also provides a more nuanced understanding of the biases present in various texts. The main contributions of this paper are:
We generate a corpus containing biases in varied domains, annotated by a team of LLM agents.
We train an NLP model for multi-label named-entity recognition, enhancing bias detection specificity and insight. To the best of our knowledge, this is the first work to provides token-level multi-label bias detection.
@@ 2 Related Works
The detection of bias in natural language processing (NLP) is a critical area of research, particularly given the increasing use of large language models (LLMs) across various domains [7, 9, 33, 18, 8, 29, 30]. Traditional techniques for bias detection often rely on human annotators to label datasets. While this approach has been instrumental in creating foundational resources, it is inherently limited by the annotators’ subjective perspectives, which can introduce their own biases into the annotation process [25, 26]. This limitation often leads to a narrow understanding of bias, especially in regard to implicit biases that are subtle and context-dependent [13, 14, 28].
@@ 2.2 Bias Detection
3 Methodology
@@
Religious | [christians, muslims, jewish people, hindus, buddhists, sikhs, atheists, agnostics | ||
@@
3.1 Dataset Generation
With modern synthetic training data labeling techniques, we can create a comprehensive dataset encapsulating our novel entities while avoiding the labor-intensive and potentially subjective human annotation process [34, 31, 2]. In addition to a synthetic data annotation pipeline, we also use a language model to synthetically generate the underlying corpus for better coverage and structural consistency of the entities we aim to classify [17, 20, 11]. The overall dataset generation pipeline can be found in Figure 1.
3.1.1 Corpus Creation
Since no existing dataset includes the specific entities of generalizations, unfairness, and stereotypes, we created a synthetic corpus of statements and questions. To ensure coverage across multiple domains, we developed four lists to utilize in prompting: Type of bias, target group, statement type, and sentiment, as shown in Table 1.
Using Mistral-7B [15], selected for its lack of guardrails, we generated prompts by combining values from each list. This process was applied with different prompt templates and types of statements for both biased and fair statements. The prompt template used for biased statement generation is illustrated in Figure 1(a).
For the generation of 1,294 fair statements, the Sentiment was modified to indicate either “slightly positive yet fair” or “slightly negative yet fair.” Responses were formatted in JSON for easier parsing and storage. While an authentic corpus could be used for annotation and training, the synthetic corpus created offers two key advantages: a broader scope that balances domains and a rich density of targeted parts of speech.
@@ 3.1.1 Corpus Creation
Generalizations | Any broad generalization of a group or ubiquitous classifiers, including adjectives/descriptors. | B-GEN, I-GEN |
@@
3.1.2 Data Annotation
The data annotation process involves several systematic steps to ensure accuracy and consistency in labeling the generated sentences. We utilized GPT-4o and the Stanford DSPy framework [19] to annotate the generated sentences with entity labels, following methodologies in recent literature for annotation with agents [24]. As shown in Figure 1(b), the annotation can be summarized in the following steps:
Preparation of annotation: We recompiled a DSPy agent for each entity type (e.g., generalization, unfairness, stereotype) as an agent to streamline the annotation process. Each agent contains the entity definition and four examples of correct annotations. The prompts sent to an agent include the definition of the target entity along with the curated examples. Using few-shot learning helps the agent understand the context and provides guidance for accurate labeling [5, 4, 6, 16]. Their definitions are provided in Table 2.
@@ 3.1.2 Data Annotation
Annotation by agents: Each generated sentence is processed by an agent supported by a language model (LLM) for a single type of entity. We included a Suggestions feature to ensure correct word/label alignment, that allows for backtracking and corrections. The agent evaluates the sentence for the presence of the specified entity and assigns appropriate labels. The agent produces a list of NER tags for each sentence, indicating the presence or absence of the entity types.
@@
Summarizing module: After annotating for one entity type, we aggregate the labels into a comprehensive two-dimensional list. Each sub-list contains one or multiple tags corresponding to each token in the sentence.
Final compilation: The individual entity labels from each annotation are systematically combined to create a final annotated dataset, with each token categorized into one or multiple semantic part-of-speech categories using B/I/O (Beginning, Inside, Outside) labels, as shown in Table 2.
In total, we annotated 3,739 sentences, each labeled for multi-label token classification training. In Figure 2 we depict the proportions of each type of bias represented in our dataset, as labeled post-generation by gpt-4o to identify cases where more than one type of bias is represented by a sentence. In Figure 3, we depict the distribution of labels in the annotated GUS dataset. It is important to note that each token can be classified with more than one label, so the sum of all labels is greater than the total number of tokens in the GUS dataset (69,679 tokens). The GUS dataset is 54.7% statements, and 45.3% questions.
3.2 Proposed Model
To efficiently and accurately identify social biases in text, we propose a multi-label token classification model. As shown in Figure 4, we fine-tune the pre-trained model bert-base-uncased [10] for multi-label classification [23].
Rather than implementing a single entity to capture all definitions and nuances of “bias,” our model achieves more granular and accurate insights with entities chosen for their individual semantic clarity and collective comprehensiveness of social bias. By utilizing a multi-label B/I/O format to represent token-level annotations, the model can predict nested entities that span multiple words. For instance, stereotypes often span a full sentence beginning with a generalization, to which some unfairness is assigned.
3.2.1 Model Architecture
The GUS-Net model is a multi-label token classification system designed to identify social bias across three categories: generalizations, unfairness, and stereotypes. It outputs seven labels, allowing the model to capture the nuanced structure of biased language. These labels facilitate the identification of individual bias categories, and provide flexibility for overlapping and nested biases.
Input Processing.
We tokenize all sentences using the pre-trained BERT tokenizer, ensuring that token splits (such as sub-words) inherit the correct entity labels from the parent word. Each text sequence is padded to a maximum length of 128 tokens to ensure consistent input size. Since sentences are rarely longer than 128 tokens, we reduced the BERT input size from the default 512 tokens, representing a 16x reduction in self-attention elements to be processed. Correspondingly, the NER tags were converted into a dimensional vector, where each of the seven elements represents a binary label (0 or 1) for the respective entity type. These vectors were padded with values up to the full sequence length of 128 tokens, with the values being ignored during the loss calculation.
Model Fine-Tuning.
We fine-tune the pre-trained transformer model, specifically bert-base-uncased, due to its ability to capture deep contextual relationships between words, which is crucial for identifying implicit biases [10]. BERT’s bidirectional nature allows it to process the entire input sequence, ensuring that each token is evaluated in the context of its surrounding words. This feature is particularly valuable in detecting subtle and complex forms of social bias.
The model is implemented using the Hugging Face transformers library. Input text is tokenized with the pre-trained BERT tokenizer, which breaks down sentences into sub-word units while preserving word boundaries. Each token sequence is padded to a length of 128 tokens, and the corresponding labels are mapped to ensure proper alignment. By reducing the input size from the default 512 tokens to 128, we optimize the model’s computational efficiency without sacrificing performance for typical sentence lengths in the dataset.
3.2.2 Loss Function
Given the significant class imbalance in our dataset, where certain entities like STEREO are underrepresented compared to frequent entities like O, we employed focal loss to address this challenge [22]. Standard binary cross-entropy (BCE) tends to focus on the majority class, leading to poor performance on rare classes. In this paper, we use focal loss calculated over all tokens, defined as:
where is the predicted probability for the true class; reduces the impact of well-classified examples, helping the model prioritize rare or difficult examples; balances the contribution of positive and negative samples, ensuring underrepresented entities receive more focus; down-weights well-classified examples, allowing the model to concentrate on harder-to-predict instances.
4 Experiments
4.1 Settings
Task Description and Metrics.
In this paper, we performed a token-level multi-label classification task, where each sentence is annotated with one or more labels to facilitate the identification of biased entities across token sequences. Token-level classification is essential for bias detection because biases can often be nuanced and context-dependent. By focusing on individual tokens, we can capture subtle implications and associations that may be overlooked in a sentence-level analysis. Moreover, we opted for multi-label classification instead of multi-class classification to better reflect the complex nature of biases, that may fall into one or many entity classes. A single-class approach, like the one proposed for Nbias, would limit our ability to capture the diversity of biases present in the text, as a statement may contain multiple biases simultaneously.
Metrics.
To evaluate the model’s performance, we utilized a variety of metrics to assess its ability to accurately identify biased entities:
Hamming Loss: This metric measures the fraction of incorrect labels over all tokens in the sequence, accounting for multi-label classification. It is defined as:
where is the total number of tokens, is the true label for the -th token, is the predicted label, and is an indicator function that evaluates whether the true label differs from the prediction.

### v3

# The GUS Framework: Benchmarking Social Bias Classification with Discriminative (Encoder-Only) and Generative (Decoder-Only) Language Models

@@ Abstract
The detection of social bias in text is a critical challenge, particularly due to the limitations of binary classification methods. These methods often oversimplify nuanced biases, leading to high emotional impact when content is misclassified as either “biased” or “fair.” To address these shortcomings, we propose a more nuanced framework that focuses on three key linguistic components underlying social bias: Generalizations, Unfairness, and Stereotypes (the GUS framework). The GUS framework employs a semi-automated approach to create a comprehensive synthetic dataset, which is then verified by humans to maintain ethical standards. This dataset enables robust multi-label token classification. Our methodology, which combines discriminative (encoder-only) models and generative (auto-regressive large language models), identifies biased entities in text. Through extensive experiments, we demonstrate that encoder-only models are effective for this complex task, often outperforming state-of-the-art methods, both in terms of macro and entity-wise F1-score and Hamming loss. These findings can guide the choice of model for different use cases, highlighting the GUS framework’s effectiveness in capturing explicit and implicit biases across diverse contexts, and offering a pathway for future research and applications in various fields. GUS resources can be found here: https://huggingface.co/collections/ethical-spectacle/gus-net-social-bias-ner-66edfe93801ea45d7a26a10f.
[…]
@@ 1 Introduction
The importance of social bias analysis in natural language processing (NLP) is increasing [1], particularly as communication increasingly relies on Large Language Models (LLMs) [2] across various domains such as education [3] and business [4]. Social bias can influence public perception and decision-making, often subtly reinforcing stereotypes or discriminatory practices. While explicit bias, which refers to overt prejudice or favoritism, is relatively easy to identify, implicit bias involves more subtle and often unconscious associations or attitudes[5]. Detecting and mitigating implicit bias in the text is significantly more challenging, as perceptions of bias can vary greatly depending on the context, including the perspectives of readers and speakers.
@@
For example, consider the phrase, ”Hard-working immigrants contribute significantly to society.” To some, this statement may appear positive, acknowledging the effort and diligence of immigrants. However, from another perspective, it might be perceived as implicitly biased, suggesting that immigrants are expected to work harder than others to be valued or accepted. This subtle implication can reinforce stereotypes that separate immigrants from native citizens, placing an undue burden of proof on their worthiness. Such subjectivity highlights the complexity of implicit bias detection, making it a critical area of research within NLP [6, 7, 8].
In state-of-the-art research, much of the focus remains on detecting biases at the sentence level [9, 10, 11]. While this approach is useful, there is a growing need for more granular bias detection. Token classification [12], or named-entity recognition (NER), enables the identification of specific words or tokens contributing to biased sentiment. This approach is often more suitable than sequence classification, which involves a higher level of abstraction that can lead to ambiguity and disagreement. However, research on linguistic bias detection at the granular level remains limited. For instance, Nbias [8] focuses on a single entity like ”BIAS,” but broader and more comprehensive frameworks are still lacking. This can be attributed to the resource-intensive nature of annotating and reviewing input text with token classes. For example, finding agreements between annotators is practical when input text is assigned a single label, as in binary classification. However, enforcing consistency of entity usage and boundaries is exponentially more difficult in human token-labeling due to personal tendencies and a less controlled environment than can now be created with LLM agents for labeling.
Modern named-entity recognition methods vary, often specifically designed for the environment they’ll be deployed. For example, social media posts tend to be unstructured or grammatically nuanced, incorporating pragmatics that diverge from conventional natural language. To handle this, recent methods incorporate data augmentation, semantic transformation, and multiple models/architectures such as BERT and LSTM to build a robust framework [13, 14, 15]. However, named-entity recognition training methods have also shown susceptibility to bias in the training data, often underperforming on specific demographics across datasets. Further, de-biasing of training data isn’t likely to improve the performance on under-represented demographics, instead fair corpus distribution is imperative [16].
One of the primary challenges in this domain is the scarcity of suitable datasets for training token classification models. While implicit bias can manifest in subtle forms, such as word choice, narrative framing, or the omission of certain viewpoints, capturing these patterns in a dataset presents unique challenges. Existing datasets, such as the Media Bias Annotation Dataset (MBIC) [17] and Bias Annotations By Experts (BABE) [18], rely on human annotators. However, the annotation process varies in sophistication. For example, MBIC outsources annotations to non-experts via Amazon Mechanical Turk, while BABE employs in-house expert annotators and ensures consensus among them. Although useful, both approaches come with trade-offs in terms of cost, dataset size, or label quality, which are further exacerbated as task complexity and granularity increase. As a result, it has become conventional to automate the extrapolation of human-labeled data, incorporating humans-in-the-loop for validation.
In response to these challenges, this paper introduces the Generalizations, Unfairness, and Stereotypes Dataset (GUS Dataset), which captures three types of bias commonly discussed in legal and psychological literature [19, 20, 21, 22]: Generalizations, Unfairness, and Stereotypes. The GUS Dataset leverages generative AI and automated agents to construct an optimal dataset for bias detection. This approach improves upon traditional methods by combining LLM reasoning with human-in-the-loop supervision, resulting in more accurate, granular, and comprehensive social bias patterns and labels across domains.
The main contributions of this paper are as follows:
1.
We generate a corpus of 3,739 text snippets depicting biases across varied domains such as religious bias, racial bias, political bias, and others. The snippets are labeled through an AI-human collaborative process, where a team of LLM agents first identifies biased words up to 4/5-grams, followed by a review by five human annotators to verify the accuracy of the labels.
2.
We format the annotations into a list of lists where each input token can be assigned multiple classes, for a multi-label token classification problem. To the best of our knowledge, this is the first work to implement data and methods for multi-label token classification with entities such as Generalizations (B-GEN, I-GEN), Unfairness (B-UNFAIR, I-UNFAIR), Stereotypes (B-STEREO, I-STEREO), or Neutral (O).
3.
We benchmark two families of models: encoder-only models (e.g., BERT-like architectures) and decoder-only auto-regressive models (e.g., LLMs), for multi-label named-entity recognition, trained on the GUS Dataset. This is the first work to provide token-level multi-label bias detection.
4.
@@
Empirical results on two model families show that encoder-only models are more effective at identifying nuanced biases, often outperforming generative models in terms of accuracy, F1-score, and Hamming loss, while maintaining computational efficiency. However, further investigation is needed to assess the reasoning capabilities of these models, as this aspect was not explicitly evaluated in our study.
@@ 2 Related Works
The detection of social bias in natural language processing (NLP) is a critical area of research, particularly given the increasing use of large language models (LLMs) across various domains [23, 24, 2, 3, 25, 4, 26]. Traditional techniques for bias detection often rely on human annotators to label datasets. While this approach has been instrumental in creating foundational resources, it is inherently limited by the annotators’ resources and expertise. This limitation often leads to datasets/frameworks that contain a narrow understanding of bias, especially in regards to implicit biases that are subtle and context-dependent [1, 6, 7].
@@ 2.2 Bias Detection
3 GUS-Net Framework for Social Bias Detection
Problem Definition We aim to detect and classify words and phrases in text that express social bias—specifically generalizations, unfair language, and stereotypes—using a named entity recognition approach with multi-label classification. This is a token-level multi-label classification task, where each sentence is annotated with one or more labels to facilitate the identification of biased entities across token sequences.
3.1 Data Preparation
Team Formation
The annotation team consisted of five members, each with a background in computational linguistics or social sciences. All team members held at least a bachelor’s degree in science, ensuring a strong foundation in both technical and domain-specific knowledge. Guidelines were provided to ensure consistency in the annotation process, including detailed instructions on entity definitions, labeling conventions, and conflict resolution protocols. Regular meetings were held to address ambiguities and maintain alignment across the team.
Synthetic Data
To construct a comprehensive dataset for social bias classification, we leverage modern synthetic training data techniques to avoid the labor-intensive and potentially subjective human annotation process [28, 29, 30]. Our approach consists of two main steps: synthetic annotation and corpus generation using a language model for enhanced coverage and structural consistency [31, 32, 33]. Figure 2 illustrates the dataset generation pipeline, and Table 1 shows the arguments combined to form prompts across domains.
Human-in-the-Loop Review and Pre-processing (Corpus)
Before annotation, the synthetic corpus underwent qualitative review by human evaluators. Reviewers examined 5-10 samples per bias type and demographic group, ensuring that generated statements were realistic and free of counterfactual stereotypes. Reviewers were also allowed to remove unrealistic cases or refine prompt categories where necessary. However, such instances were rare, confirming that Mistral-7B effectively generated relevant bias-based statements. Finally, a quantitative distribution analysis of the prompt arguments ensured balanced representation across bias types and demographic groups. Once approved, the dataset proceeded to the annotation stage.
@@ 3.1 Data Preparation
| Religious | [christians, muslims, jewish people, hindus, buddhists, sikhs, atheists, agnostics, …] | ||
@@
Data Annotation
The data annotation process involves several systematic steps to ensure accuracy and consistency in labeling the generated sentences. We utilized GPT-4o and the Stanford DSPy framework [34] to annotate the generated sentences with entity labels, following methodologies in recent literature for annotation with agents [35]. As shown in Figure 2, the annotation can be summarized in the following steps:
Preparation of annotation: We recompiled a DSPy agent for each entity type (e.g., generalization, unfairness, stereotype) as an agent to streamline the annotation process. Each agent contains the entity definition and four examples of correct annotations. The prompts sent to an agent include the definition of the target entity along with the curated examples. Using few-shot prompting helps the agent understand the context and provides guidance for accurate labeling [36, 37, 38, 39]. Their definitions are provided in Table 2.
@@
| Generalizations | Any broad generalization of a group or ubiquitous classifiers, including adjectives and descriptors. | B-GEN, I-GEN |
@@
Annotation by agents: Each generated sentence is processed by an agent supported by an LLM for a single type of entity. We included a Suggestions feature to ensure correct word and label alignment, which allows for backtracking and corrections. The agent evaluates the sentence for the presence of the specified entity and assigns appropriate labels in a single-label Beginning/Inside/Outside (BIO) format [40]. Each agent produces a list of named-entity recognition (NER) tags for each sentence, indicating the presence or absence of the entity types.
@@
Summarizing module: After annotating a sentence for each entity type independently, the labels were systematically aggregated into a comprehensive two-dimensional list. Each sub-list contains one or multiple tags for each word in the text, as shown in Figure 2. At this stage, spaces were used as delimiters for checking the alignment of labels, so the labels are word-level, not token-level.
Human-in-the-Loop Review (Annotations)
In total, 3,739 sentences were annotated, each labeled for multi-label token classification. Figure 3(a) depicts the distribution of each type of bias represented in the GUS Dataset, as labeled post-generation by GPT-4o. This was done as part of our quantitative review process, and re-labeling type-of-bias post-generation was done to capture cases where a sentence represents more than one type of bias. While the distribution of biases in the dataset was generally balanced, the distribution of entities was not. Figure 3(b) depicts the distribution of token labels in the annotated GUS dataset. Since each token can be classified with more than one label, the total number of labels is greater than the total number of tokens in the GUS dataset (69,679 tokens). The unbalanced distribution of token labels is inherent to the task, as some entities like UNFAIR are less common than others like O (neutral), and was considered while constructing the training architecture for benchmarking. The GUS dataset is 54.7% statements and 45.3% questions. Having satisfied the human reviewers’ qualitative and quantitative checks, the annotated dataset was deemed ready for downstream tasks, including model training and evaluation.
3.2 Multi-Label Token Classification for Social Bias Detection
Given an input text sequence , where each token represents a word or subword in the sequence, the goal is to classify each token into one or more categories of social bias:
1.
Generalization (GEN)
2.
Unfairness (UNFAIR)
3.
Stereotype (STEREO)
Since tokens can belong to multiple categories simultaneously, this constitutes a multi-label token classification problem. To preserve sequence context and detect nested biases, we employ a B/I/O (Beginning, Inside, Outside) tagging scheme, where:
1.
B-ENTITY: Marks the beginning of a bias-related entity.
2.
I-ENTITY: Marks the continuation of the entity.
3.
O (Outside): Marks tokens that do not belong to any bias category.
We define a multi-hot label matrix for the sequence:
where:
1.
is the number of tokens in the input sequence.
2.
(corresponding to GEN, UNFAIR, and STEREO).
For each token , the model predicts a label vector:
where:
The function maps the input sequence to a predicted label matrix :
To train the model, we minimize the binary cross-entropy (BCE) loss across all tokens and bias categories:
To mitigate class imbalance, we apply focal loss:
The model learns to assign multiple bias labels per token while handling class imbalances effectively using focal loss. The function can be instantiated as a transformer-based model fine-tuned on the GUS dataset.
3.3 Benchmarking Discriminative and Generative Models
In our experiments, we benchmark two distinct classes of models for the multi-label token classification task: discriminative (encoder-only) models and generative (decoder-only) models.
Discriminative Models (Encoder-Only)
For the discriminative approach, we fine-tune pre-trained encoder-only language models. These models extract token embeddings and add a dense layer to map each token to the three target bias categories. The input processing pipeline involves truncating or padding sequences to a fixed length of 128 tokens, aligning word-level labels with subword tokens (with subword tokens inheriting the label of their parent word), and converting entity tags into a matrix. To address class imbalances, we incorporate focal loss during training.
Our evaluation includes transformer-based encoders such as DistilBERT (66M parameters), BERT-base-uncased (110M parameters), and RoBERTa-base (123M parameters), all implemented using the Hugging Face transformers library.
Generative Models (Decoder-Only)
Generative, auto-regressive LLMs are adapted to perform token-level labeling through instruction fine-tuning and few-shot prompting. Due to their sequential generation process, these models are not inherently designed for token-level classification. To accommodate this, we format each input as a chatML prompt that includes a user message (providing instructions, target entities, and the input text) and a system message containing the corresponding true labels from the GUS dataset. This prompt-based format allows us to evaluate the models both quantitatively (by assessing alignment accuracy and standard metrics) and qualitatively (through their reasoning ability).
For parameter-efficient fine-tuning, we employ low-rank adaptations (LoRA) using Unsloth with a LoRA rank of 16 and an of 16; in one instance, a model is quantized to 4-bit precision due to memory constraints. Our experiments with generative models involve Llama 3.3 (70B parameters), Llama 3.2 (3B parameters), Phi 3 Medium (14B parameters), and Phi 3.5 Mini (4B parameters).
Unified Evaluation.
Both discriminative and generative models are benchmarked on the GUS dataset. While encoder-only models offer efficient, direct token-level classification, decoder-only models provide enhanced reasoning capabilities and flexible, prompt-based responses. We evaluate the models using standard token classification metrics (such as precision, recall, and F1 score) along with qualitative assessments of model reasoning.
4 Experimental Settings
4.1 Hardware Setting
All experiments for discriminative (encoder-only) models were conducted on a single NVIDIA T4 GPU with 16GB of memory, while generative auto-regressive (decoder-only) 111In the rest of this document, we use encoder-only to represent discriminative (encoder-only) models and decoder-only to represent generative auto-regressive (decoder-only) LLMs. models were fine-tuned on an A100 with 40GB of memory; both were executed on Google Colab. We also tested decoder-only models for the prompting in few-shots settings with same hardware configurations. The training was implemented using PyTorch and the transformers library, executed on Ubuntu 20.04 with Python 3.8. We used PyTorch Lightning to streamline the training loops and logging mechanisms for encoder fine-tuning, and Unsloth for parameter-efficient fine-tuning of LLMs.
4.2 Evaluation Strategy
Evaluation Data
The dataset developed and used in this study is detailed in Section 3.1. It includes 3,739 annotated samples evenly distributed across types of bias, initially labeled by GPT-4o and subsequently reviewed by human annotators. This dataset was used for fine-tuning both encoder-only and decoder-only models. The data is divided into training, validation, and test sets with an 70-15-15 split ratio by random sample. The BABE [17] dataset was used as an out of distribution corpus in the same settings as we used our dataset.
Evaluation Metrics
This work is a token level classification task so we utilized a variety of metrics that are commonly used for this task in the related works [41] to assess its ability to accurately identify biased entities. These metrics are accuracy-based for evaluating multi-label classification problems.
1.